AutoSave OFF OFF A SOCIET a Home Insert Draw Design Layout References Mailings Review View Tell me Share Comments Aalcedie Aamake a claim based on the presented data in the table, then answer set questions about the claim.
Answer #1

Make a claim: The average food affordability of States is more than 9.

Level of significance: g= 0,05

Null hypothesis: H_{0}:\mu=9

Alternate hypothesis: H_{1}:\mu>9

The random variable X represents the sample mean of the food affordability.

Distribution: I shall use a t-distribution since the population variance is not knownwith n-1=14 degrees of freedom

Test statistic: X- t= s/n

t=\frac{9.2067-9}{2.4665/\sqrt{15}}

t=\frac{0.2067}{2.4665/3.8730}

t=\frac{0.2067}{0.6368}

t=0.3246

Type of test: It is an one tailed test and is right tailed.

p-value: The p=0.3752

The p-value for this test means that around 38 in 100 times, we could have concluded that the population mean food affordability is more than 9 and in fact the true availability 9.

---------------------------------------------------------------------------------------------------------------------------------------------------------------

For calculating the mean and variance, we make use of the table of sum and sum of squares.

State Food affordability x^2
1 7.8 60.84
2 9.9 98.01
3 6.2 38.44
4 11.2 125.44
5 8.3 68.89
6 7.8 60.84
7 6 36
8 9.7 94.09
9 6 36
10 15.5 240.25
11 10.2 104.04
12 9.9 98.01
13 9.4 88.36
14 8.8 77.44
15 11.4 129.96
Total 138.1 1356.61

Mean of x, =1 TS ci η

\bar{x}=\frac{138.1}{15}=9.2067

Sample variance of x, s^{2}=\frac{1}{n-1}\left [ \sum x_{i}^{2}-\frac{1}{n}\left ( \sum x_{i} \right )^{2} \right ]

  s^{2}=\frac{1}{15-1}\left [ 1356.61-\frac{1}{15}\left (138.1 \right )^{2} \right ]

  s^{2}=\frac{1}{14}\left [ 1356.61-\frac{19071.61}{15} \right ]

  s^{2}=\frac{1}{14}\left [ 1356.61-1271.4407 \right ]

s^{2}=\frac{1}{14}\left [ 85.1693 \right ]

  s^{2}=6.0835

  s=2.4665